# The Spacetime of Diffusion Models: An Information Geometry Perspective

### v2

@@ Abstract
We present a novel geometric perspective on the latent space of diffusion models. We first show that the standard pullback approach, utilizing the deterministic probability flow ODE decoder, is fundamentally flawed. It provably forces geodesics to decode as straight segments in data space, effectively ignoring any intrinsic data geometry beyond the ambient Euclidean space. Complementing this view, diffusion also admits a stochastic decoder via the reverse SDE, which enables an information geometric treatment with the Fisher-Rao metric. However, a choice of ${\bm{x}}_{T}$ as the latent representation collapses this metric due to memorylessness. We address this by introducing a latent spacetime ${\bm{z}}=({\bm{x}}_{t},t)$ that indexes the family of denoising distributions $p({\bm{x}}_{0}|{\bm{x}}_{t})$ across all noise scales, yielding a nontrivial geometric structure. We prove these distributions form an exponential family and derive simulation-free estimators for curve lengths, enabling efficient geodesic computation. The resulting structure induces a principled Diffusion Edit Distance, where geodesics trace minimal sequences of noise and denoise edits between data. We also demonstrate benefits for transition path sampling in molecular systems, including constrained variants such as low-variance transitions and region avoidance. Code is available at: https://github.com/rafalkarczewski/spacetime-geometry.
@@ 1 Introduction
In generative models, a common way to study the intrinsic geometry of the data is to pull back the ambient (Euclidean) metric onto the latent space (Arvanitidis et al., 2018, 2022). Equipped with this pullback metric, shortest paths (i.e., geodesics) in the latent space decode to realistic transitions along data that lie on a lower-dimensional submanifold.
@@
We then turn our attention to the decoding distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$ given by the reverse SDE. We propose an alternative Fisher-Rao geometry, which measures how the denoising distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$ changes when manipulating the latent ${\bm{x}}_{t}$. We introduce the Fisher-Rao metric $\mathbf{G}({\bm{x}}_{t},t)$ that varies with both state and time over the latent spacetime $({\bm{x}}_{t},t)$ (Fig.˜1).
@@ 6.1 Sampling trajectories
We compare the trajectories obtained by solving the PF-ODE ${\bm{x}}_{0}({\bm{x}}_{T})$ (Eq. 3) with geodesics between the same endpoints ${\bm{x}}_{0},{\bm{x}}_{T}$. For a toy example of 1D mixture of Gaussians, we observe the geodesics curving less than the PF-ODE trajectories in the early sampling (high $t$), while being indistinguishable for lower values of $t$ (See Fig.˜3 left and Section˜F.1 for details).
We find only marginal perceptual difference between the PF-ODE sampling trajectories and the geodesics in the EDM2 ImageNet-512 model (Karras et al., 2024). The geodesic appears to generate information slightly earlier, but the difference is minor (See Fig.˜3 right, and Section˜F.2 for details).
@@ 6.2 Diffusion Edit Distance
where $\bm{\gamma}$} is the spacetime geodesic between $({\bm{x}}^{a},0)$ and $({\bm{x}}^{b},0)$. For numerical stability, we anchor endpoints at a small $t_{\min}>0$ rather than at 0.
@@ 6.3 Transition path sampling
To construct a transition path between two low-energy states ${\bm{x}}_{0}^{1}$ and ${\bm{x}}_{0}^{2}$, we estimate the spacetime geodesic $\bm{\gamma}$} between them using a denoiser model $\hat{{\bm{x}}}_{0}({{\bm{x}}_{t}})\approx\mathbb{E}[{\bm{x}}_{0}|{{\bm{x}}_{t}}]$ with Proposition˜5.1, as shown in Fig.˜5. At each interpolation point $s\in[0,1]$, the geodesic defines a denoising Boltzmann distribution $p({\bm{x}}|{\bm{\gamma}}_{s})$ where $U({\bm{x}}|{\bm{\gamma}}_{s})$ is the energy at that spacetime location. See Section˜F.3 for details.
@@
Table 1: Spacetime geodesics outperform methods tailored to transition path sampling. Parentheses denote extra energy evaluations used to generate training data for the base diffusion model, which do not scale with the number of generated paths. Baseline details in Appendix˜G.
@@
Spacetime geodesic (Ours) 37.66 ± 0.61 128K (+16M)
@@
We compute a spacetime geodesic connecting two molecular configurations of Alanine Dipeptide, as in Holdijk et al. (2023). In Fig.˜5, the energy landscape is visualized over the dihedral angle space, with a neural network used to approximate the potential energy $U$. Using our trained denoiser $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, we estimate the expectation parameter $\bm{\mu}$}, which allows us to compute and visualize a geodesic trajectory through spacetime. Transition paths were generated using Algorithm˜1. See Section˜F.3 for details.
@@
We considered Holdijk et al. (2023); Du et al. (2024); Raja et al. (2025) and adopt Doob’s Lagrangian (Du et al., 2024); the others were excluded due to reproducibility issues (see Appendix˜G). We also evaluate two MCMC two-way shooting variants (Brotzakis and Bolhuis, 2016)-uniform point selection with variable or fixed trajectory length-using transition paths from the official Du et al. (2024) code release. For each method we generate 1,000 paths and report mean MaxEnergy (lower is better) and its numerical lower bound $\min_{{\bm{\gamma}}}\max_{s}U({\bm{\gamma}}_{s})$, along with the number of energy evaluations needed for 1,000 paths. To train a base diffusion model for our method, we generated data using Langevin dynamics (16M11116M is the number of energy function evaluations to generate the training set with Langevin dynamics for the base diffusion model. We did not tune this number, and fewer evaluations may yield comparable performance. energy evaluations), a one-time cost that does not scale with the number of generated transition paths.
@@ Results.
We show in Table˜1 that our method outperforms the baselines in the MaxEnergy obtained along the transition paths. It is also considerably closer to the lower bound than to the next best baseline (MCMC-fixed length) while requiring several orders of magnitude fewer energy function evaluations. In Fig.˜6, we show a qualitative comparison of transition paths generated with our method and the baselines. Our proposed method shows improved efficiency in avoiding high-energy regions compared to MCMC. In contrast, the Doob’s Lagrangian method converged to a suboptimal solution, producing nearly identical transition paths. We discuss this in more detail in Appendix˜G.
@@ 8 Limitations
Although our framework defines geodesics between any noisy samples, optimizing between nearly clean ones is numerically unstable because their denoising distributions collapse to Dirac deltas, making Fisher-Rao (via local KL) distances effectively infinite. Therefore, consistent with diffusion practice (Song et al., 2021; Lu et al., 2022), we choose endpoints with non-negligible noise for tractable optimization (details in Appendix˜F).
The proposed distance metric DiffED (Section˜6.2) is considerably slower (details in Section˜F.2) than established image similarity metrics such as LPIPS (Zhang et al., 2018), or SSIM (Wang et al., 2004). Exploring a distillation strategy involving training a separate model trained to predict DiffED is a possible future research direction.
@@ C.1 Information geometry in exponential families
$\mathcal{E}_{\mathrm{IG}}({\bm{\gamma}})=\frac{1}{2}\int_{0}^{1}\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right)ds.$ (42)
@@
$\mathcal{E}_{\mathrm{IG}}({\bm{\gamma}})$ $=\frac{1}{2}\int_{0}^{1}\|\dot{{\bm{\gamma}}}_{s}\|_{\mathbf{G}_{\mathrm{IG}}}^{2}ds=\frac{1}{2}\int_{0}^{1}\dot{{\bm{\gamma}}}_{s}^{\top}\mathbf{G}_{\mathrm{IG}}({{\bm{\gamma}}_{s}})\dot{{\bm{\gamma}}}_{s}ds$
$=\frac{1}{2}\int_{0}^{1}\dot{{\bm{\gamma}}}_{s}^{\top}\left(\frac{\partial{\bm{\eta}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\right)^{\top}\left(\frac{\partial{\bm{\mu}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\right)\dot{{\bm{\gamma}}}_{s}ds$
$=\frac{1}{2}\int_{0}^{1}\left(\frac{\partial{\bm{\eta}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\dot{{\bm{\gamma}}}_{s}\right)^{\top}\left(\frac{\partial{\bm{\mu}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\dot{{\bm{\gamma}}}_{s}\right)ds$
$=\frac{1}{2}\int_{0}^{1}\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right)ds.$
@@ Appendix D Kullback-Leibler divergence in exponential families
Appendix E Qualitative example of DiffED

### v3

@@ Abstract
We present a novel geometric perspective on the latent space of diffusion models. We first show that the standard pullback approach, utilizing the deterministic probability flow ODE decoder, is fundamentally flawed. It provably forces geodesics to decode as straight segments in data space, effectively ignoring any intrinsic data geometry beyond the ambient Euclidean space. Complementing this view, diffusion also admits a stochastic decoder via the reverse SDE, which enables an information geometric treatment with the Fisher-Rao metric. However, a choice of ${\bm{x}}_{T}$ as the latent representation collapses this metric due to memorylessness. We address this by introducing a latent spacetime ${\bm{z}}=({\bm{x}}_{t},t)$ that indexes the family of denoising distributions $p({\bm{x}}_{0}|{\bm{x}}_{t})$ across all noise scales, yielding a nontrivial geometric structure. We prove these distributions form an exponential family and derive simulation-free estimators for curve lengths, enabling efficient geodesic computation. The resulting structure induces a principled Diffusion Edit Distance, where geodesics trace minimal sequences of noise and denoise edits between data. We also demonstrate benefits for transition path sampling in molecular systems, including constrained variants such as low-variance transitions and region avoidance. Code is available at https://github.com/rafalkarczewski/spacetime-geometry.
@@ 1 Introduction
In generative models, a common way to study the intrinsic geometry of the data is to pull back the ambient (Euclidean) metric onto the latent space (Arvanitidis et al., 2018; 2022). Equipped with this pullback metric, shortest paths (i.e., geodesics) in the latent space decode to realistic transitions along data that lie on a lower-dimensional submanifold.
@@
We then turn our attention to the denoising posterior distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$ given by the reverse SDE. We propose an alternative Fisher-Rao geometry, which measures how the denoising distribution $p({\bm{x}}_{0}|{\bm{x}}_{t})$ changes when manipulating the latent ${\bm{x}}_{t}$. We introduce the Fisher-Rao metric $\mathbf{G}({\bm{x}}_{t},t)$ that varies with both state and time over the latent spacetime $({\bm{x}}_{t},t)$ (Fig.˜1).
@@ 6.1 Sampling trajectories
We compare the trajectories obtained by solving the PF-ODE ${\bm{x}}_{0}({\bm{x}}_{T})$ (Eq. 3) with geodesics between the same endpoints ${\bm{x}}_{0},{\bm{x}}_{T}$. For a toy example of 1D mixture of Gaussians, we observe the geodesics curving less than the PF-ODE trajectories in the early sampling (high $t$), while being indistinguishable for lower values of $t$ (See Fig.˜3 left and Section˜G.1 for details).
We find only marginal perceptual difference between the PF-ODE sampling trajectories and the geodesics in the EDM2 ImageNet-512 model (Karras et al., 2024). The geodesic appears to generate information slightly earlier, but the difference is minor (See Fig.˜3 right, and Section˜G.2 for details).
@@ 6.2 Diffusion Edit Distance
where $\bm{\gamma}$} is the spacetime geodesic between $({\bm{x}}^{a},0)$ and $({\bm{x}}^{b},0)$. For numerical stability, we anchor endpoints at a small $t_{\min}>0$ rather than at 0. See Algorithm˜4 for DiffED pseudocode.
@@ 6.3 Transition path sampling
To construct a transition path between two low-energy states ${\bm{x}}_{0}^{1}$ and ${\bm{x}}_{0}^{2}$, we estimate the spacetime geodesic $\bm{\gamma}$} between them using a denoiser model $\hat{{\bm{x}}}_{0}({{\bm{x}}_{t}})\approx\mathbb{E}[{\bm{x}}_{0}|{{\bm{x}}_{t}}]$ with Proposition˜5.1, as shown in Fig.˜5. At each interpolation point $s\in[0,1]$, the geodesic defines a denoising Boltzmann distribution $p({\bm{x}}|{\bm{\gamma}}_{s})$ where $U({\bm{x}}|{\bm{\gamma}}_{s})$ is the energy at that spacetime location. See Section˜G.3 for details.
@@
Table 1: Spacetime geodesics outperform methods tailored to transition path sampling. Parentheses denote extra energy evaluations used to generate training data for the base diffusion model, which do not scale with the number of generated paths. Baseline details in Appendix˜H.
@@
Spacetime geodesic (Ours) 37.36 ± 0.60 16M (+16M)
@@
We compute a spacetime geodesic connecting two molecular configurations of Alanine Dipeptide, as in Holdijk et al. (2023). In Fig.˜5, the energy landscape is visualized over the dihedral angle space, with a neural network used to approximate the potential energy $U$. Using our trained denoiser $\hat{{\bm{x}}}_{0}({\bm{x}}_{t})$, we estimate the expectation parameter $\bm{\mu}$}, which allows us to compute and visualize a geodesic trajectory through spacetime. Transition paths were generated using Algorithm˜1. See Section˜G.3 for details.
@@
We considered Holdijk et al. (2023); Du et al. (2024); Raja et al. (2025) and adopt Doob’s Lagrangian (Du et al., 2024); the others were excluded due to reproducibility issues (see Appendix˜H). We also evaluate two MCMC two-way shooting variants (Brotzakis and Bolhuis, 2016)-uniform point selection with variable or fixed trajectory length-using transition paths from the official Du et al. (2024) code release. For each method we generate 1,000 paths and report mean MaxEnergy (lower is better) and its numerical lower bound $\min_{{\bm{\gamma}}}\max_{s}U({\bm{\gamma}}_{s})$, along with the number of energy evaluations needed for 1,000 paths. To train a base diffusion model for our method, we generated data using Langevin dynamics (16M111+16M is the number of energy function evaluations to generate the training set with Langevin dynamics for the base diffusion model. We did not tune this number, and fewer evaluations may yield comparable performance. energy evaluations), a one-time cost that does not scale with the number of generated transition paths.
@@ Results.
We show in Table˜1 that our method outperforms the baselines in the MaxEnergy obtained along the transition paths. It is also considerably closer to the lower bound than to the next best baseline (MCMC-fixed length) while requiring several orders of magnitude fewer energy function evaluations. In Fig.˜6, we show a qualitative comparison of transition paths generated with our method and the baselines. Our proposed method shows improved efficiency in avoiding high-energy regions compared to MCMC. In contrast, the Doob’s Lagrangian method converged to a suboptimal solution, producing nearly identical transition paths. We discuss this in more detail in Appendix˜H.
@@ 8 Limitations
Although our framework defines geodesics between any noisy samples, optimizing between nearly clean ones is numerically unstable because their denoising distributions collapse to Dirac deltas, making Fisher-Rao (via local KL) distances effectively infinite. Therefore, consistent with diffusion practice (Song et al., 2021; Lu et al., 2022), we choose endpoints with non-negligible noise for tractable optimization (details in Appendix˜G).
The proposed distance metric DiffED (Section˜6.2) is considerably slower (details in Section˜G.2) than established image similarity metrics such as LPIPS (Zhang et al., 2018), or SSIM (Wang et al., 2004). Exploring a distillation strategy involving training a separate model trained to predict DiffED is a possible future research direction.
@@ 9 Conclusion
Reproducibility statement
We include the source code in our submission, which allows for reproducing the results. Our claims made in the main text are proven in the appendices. Experiment details can be found in Appendix˜G.
@@ C.1 Information geometry in exponential families
$\mathcal{E}_{\mathrm{IG}}({\bm{\gamma}})$ $=\frac{1}{2}\int_{0}^{1}\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right)ds$ (42)
$\ell_{\mathrm{IG}}({\bm{\gamma}})$ $=\int_{0}^{1}\sqrt{\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right)}ds.$ (43)
For a curve discretized uniformly into $N$ points, we have $s_{n}:=\tfrac{n}{N-1}$, ${\bm{\gamma}}_{n}:={\bm{\gamma}}(s_{n})$, and ${\bm{\mu}}_{n}:={\bm{\mu}}({\bm{\gamma}}_{n})$, ${\bm{\eta}}_{n}:={\bm{\eta}}({\bm{\gamma}}_{n})$, we have
$\mathcal{E}_{\mathrm{IG}}({\bm{\gamma}})$ $\approx\frac{N-1}{2}\sum_{n=0}^{N-2}\left({\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}\right)^{\top}\left({\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}\right)$ (44)
$\ell_{\mathrm{IG}}({\bm{\gamma}})$ $\approx\sum_{n=0}^{N-2}\sqrt{\left({\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}\right)^{\top}\left({\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}\right)}$ (45)
@@
Using Eq.˜32), we have
$\|\dot{{\bm{\gamma}}}_{s}\|_{\mathbf{G}_{\mathrm{IG}}}^{2}$ $=\dot{{\bm{\gamma}}}_{s}^{\top}\mathbf{G}_{\mathrm{IG}}({{\bm{\gamma}}_{s}})\dot{{\bm{\gamma}}}_{s}=\dot{{\bm{\gamma}}}_{s}^{\top}\left(\frac{\partial{\bm{\eta}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\right)^{\top}\left(\frac{\partial{\bm{\mu}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\right)\dot{{\bm{\gamma}}}_{s}$
$=\left(\frac{\partial{\bm{\eta}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\dot{{\bm{\gamma}}}_{s}\right)^{\top}\left(\frac{\partial{\bm{\mu}}({\bm{\gamma}}_{s})}{\partial{\bm{z}}}\dot{{\bm{\gamma}}}_{s}\right)=\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right).$
Therefore
$\mathcal{E}_{\mathrm{IG}}({\bm{\gamma}})$ $=\frac{1}{2}\int_{0}^{1}\|\dot{{\bm{\gamma}}}_{s}\|_{\mathbf{G}_{\mathrm{IG}}}^{2}ds=\frac{1}{2}\int_{0}^{1}\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right)ds$
$\ell_{\mathrm{IG}}({\bm{\gamma}})$ $=\int_{0}^{1}\|\dot{{\bm{\gamma}}}_{s}\|_{\mathbf{G}_{\mathrm{IG}}}ds=\int_{0}^{1}\sqrt{\left(\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s})\right)^{\top}\left(\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s})\right)}ds.$
For a discretized curve, the reasoning is similar to the proof of Proposition A.2 in Arvanitidis et al. (2022). We have $ds=s_{n+1}-s_{n}=\tfrac{1}{N-1}$ and we can approximate $\tfrac{d}{ds}{\bm{\eta}}({\bm{\gamma}}_{s}))\big\rvert_{s=s_{n}}\approx\frac{{\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}}{ds}$ and $\tfrac{d}{ds}{\bm{\mu}}({\bm{\gamma}}_{s}))\big\rvert_{s=s_{n}}\approx\frac{{\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}}{ds}$, and
$\mathcal{E}_{\mathrm{IG}}({\bm{\gamma}})$ $\approx\frac{1}{2}\sum_{n=0}^{N-2}\left(\frac{{\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}}{ds}\right)^{\top}\left(\frac{{\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}}{ds}\right)ds=\frac{N-1}{2}\sum_{n=0}^{N-2}\left({\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}\right)^{\top}\left({\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}\right)$
$\ell_{\mathrm{IG}}({\bm{\gamma}})$ $\approx\sum_{n=0}^{N-2}\sqrt{\left(\frac{{\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}}{ds}\right)^{\top}\left(\frac{{\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}}{ds}\right)}ds=\sum_{n=0}^{N-2}\sqrt{\left({\bm{\eta}}_{n+1}-{\bm{\eta}}_{n}\right)^{\top}\left({\bm{\mu}}_{n+1}-{\bm{\mu}}_{n}\right)}$
@@ Appendix D Kullback-Leibler divergence in exponential families
Appendix E Geodesic and DiffED pseudocode
Algorithm 2 Evaluate Curve
1:Curve ${\bm{\gamma}}:[0,1]\to\mathbb{R}^{D+1}$, number of points $N_{\gamma}$
[…]
3:for $n=0$ to $N_{\gamma}-1$ do
4:  ${\bm{z}}_{n}\leftarrow{\bm{\gamma}}(s_{n})$
5:  Decompose ${\bm{z}}_{n}=({\bm{x}}_{t}^{(n)},t_{n})$
6:  ${\bm{\eta}}_{n}\leftarrow{\bm{\eta}}({\bm{x}}_{t}^{(n)},t_{n})$ using Eq.˜15
7:  ${\bm{\mu}}_{n}\leftarrow{\bm{\mu}}({\bm{x}}_{t}^{(n)},t_{n})$ using Eq.˜16
8:end for
9:return $\{{\bm{\eta}}_{n}\},\{{\bm{\mu}}_{n}\}$
Algorithm 3 Spacetime Geodesic Estimation
[…]
4:Initialize cubic spline ${\bm{\gamma}}_{{\bm{\theta}}}$ with ${\bm{\gamma}}_{{\bm{\theta}}}(0)={\bm{z}}_{a}$, ${\bm{\gamma}}_{{\bm{\theta}}}(1)={\bm{z}}_{b}$
[…]
6:  $\{{\bm{\eta}}_{n}\},\{{\bm{\mu}}_{n}\}\leftarrow\textsc{EvaluateCurve}({\bm{\gamma}}_{{\bm{\theta}}},N_{\gamma})$
7:  $\mathcal{E}({\bm{\gamma}}_{\bm{\theta}})\leftarrow\frac{N_{\gamma}-1}{2}\sum_{n}({\bm{\eta}}_{n+1}-{\bm{\eta}}_{n})^{\top}({\bm{\mu}}_{n+1}-{\bm{\mu}}_{n})$ $\triangleright$ Energy estimate Eq.˜14
8:  ${\bm{g}}\leftarrow\nabla_{{\bm{\theta}}}\mathcal{E}({\bm{\gamma}}_{\bm{\theta}})$
9:  ${\bm{\theta}}\leftarrow{\bm{\theta}}-\eta{\bm{g}}\triangleright$ Gradient descent update of curve’s parameters
10:end for
11:return ${\bm{\gamma}}_{\bm{\theta}}$
Algorithm 4 Diffusion Edit Distance Estimation
[…]
6:return $d_{\mathrm{DE}}({\bm{x}}_{a},{\bm{x}}_{b})$
Appendix F Qualitative example of DiffED